# Toward Socially-Aware LLMs: A Survey of Multimodal Approaches to Human Behavior Understanding

### v1

@@ Abstract.
LLM-powered multimodal systems are increasingly used to interpret human social behavior, yet how researchers apply the models’ ”social competence” remains poorly understood. This paper presents a systematic literature review of 176 publications across different application domains (e.g., healthcare, education, and entertainment). Using a four-dimensional coding framework (application, technical, evaluative, and ethical), we find (1) frequent use of pattern recognition and information extraction from multimodal sources, but limited support for adaptive, interactive reasoning; (2) a dominant ”modality-to-text” pipeline that privileges language over rich audiovisual cues, striping away nuanced social cues; (3) evaluation practices reliant on static benchmarks, with socially grounded, human-centered assessments rare; and (4) Ethical discussions focused mainly on legal and rights-related risks (e.g., privacy), leaving societal risks (e.g., deception) overlooked—or at best acknowledged but left unaddressed. We outline a research agenda for evaluating socially competent, ethically informed, and interaction-aware multi-modal systems.
@@ 1. Introduction
This paper addresses the gap through a systematic literature review of LLM-powered multimodal systems that are designed to understand human social behavior. Our analysis synthesizes research across technical venues in AI and applied communities such as HCI, education, and healthcare. Specifically, we examine the following four perspectives:
  • RQ1 (application): What social intelligence of LLM-powered multimodal systems is applied to human social behavior analysis across different contexts?
  • RQ2 (technical): How is LLM-powered multimodal systems social intelligence technically operationalized?
  • RQ3 (evaluative): How is the performance of socially intelligent LLM-powered multimodal systems evaluated?
  • RQ4 (ethical): What are the primary ethical challenges and risks associated with socially intelligent LLM-powered multimodal systems?
@@ 3.3. Coding Procedure
4. RQ1: Scope and Nature of Social Intelligence in LLM-powered Multimodal Systems
What social intelligence of LLM-powered multimodal systems is applied to human social behavior analysis across different contexts?
@@ 4. RQ1: Scope and Nature of Social Intelligence in LLM-powered Multimodal Systems
4.1. Social Intelligence Functions
To understand how LLM-powered multimodal systems operationalize social intelligence, we analyzed both the application domains these systems target, and the types of human behaviors they are designed to recognize and interpret.
@@ 4.1. Social Intelligence Functions
Our analysis also reveals clear differences in cue usage across application contexts (Figure. 4). For example, healthcare and well-being systems tend to integrate a broad spectrum of cues—including facial expressions, prosody, and motion—to support emotionally responsive interactions, such as in cognitive therapy. In contrast, education and training applications emphasize verbal and language cues, aligning with their focus on dialogue analysis and instruction. Security and surveillance systems skew heavily toward motion and action recognition, while domains like autonomous driving prioritize both complex goal-directed actions and spatial body movement cues.
Figure 4. Distribution of Behavioral Cue Types Across Application Domains. This heatmap illustrates the frequency of different behavioral cues analyzed within key application areas. It reveals distinct patterns of focus; for instance, ”Healthcare & Well-being” utilizes a wide range of cues, whereas ”Education & Training” predominantly focuses on verbal and linguistic cues.
@@ 4.2. Behavior Characteristics
To characterize the nature of social behaviors analyzed in LLM-powered multimodal systems, we examined three key dimensions: interaction structure (i.e., the number of interacting participants), temporal scale (i.e., the time span of behavior under analysis), and social units (i.e., the building blocks of social analysis). Together, these dimensions offer a structured view of the complexity of human behavior addressed by current systems (Figure. 5).
@@ 4.4. Operationalized Social Intelligence Competencies
Major Takeaways–Scope and Nature of Social Intelligence (RQ1) Dominance of Perception & Recognition: Nearly all of the 176 works focus on social perception (100%) and reasoning (95%), while social interaction (41.8%) and creativity (18%) are rarely implemented. Short-Term, Individual-Centric Focus: 94% of studies analyze immediate or short-term behavior and 63% focus on single-person analysis, with very limited modeling of long-term trajectories or multi-party dynamics. Uneven Cue Usage: Verbal/language cues dominate, while nuanced non-verbal signals like gaze, facial expressions, and prosody are underutilized despite their importance for social intelligence. Observer Role Prevails: Most systems act as post-hoc observers rather than active participants, highlighting a gap between analyzing and co-constructing social interaction.
5. RQ2: Technical Operationalization of Social Intelligence in Multimodal LLM Systems
How is LLM-powered multimodal systems social intelligence technically operationalized?
@@ 5.2. Operationalization Technique
6. RQ3: Evaluation of Social Intelligence in Multimodal LLM Systems
How is the performance of socially intelligent LLM-powered multimodal systems evaluated?
@@ 6.2. System Evaluation Mechanism
7. RQ4: Ethical Challenges and Risks of Socially Intelligent Multimodal LLM Systems
What are the primary ethical challenges and risks associated with socially intelligent LLM-powered multimodal systems?
7.1. Setup and Coding Approach

### v2

# The Social Gaze of LLMs: A Literature Review of Multimodal Approaches to Human Behavior Understanding

@@ Abstract.
LLM-powered multimodal systems are increasingly used to interpret human behavior, yet how researchers apply the models’ ”social competence” remains poorly understood. This paper presents a systematic literature review of 176 publications across different application domains (e.g., healthcare, education, and entertainment). Using a four-dimensional coding framework (application, technical, evaluative, and ethical), we find (1) frequent use of pattern recognition and information extraction from multimodal sources, but limited support for adaptive, interactive reasoning; (2) a dominant ”modality-to-text” pipeline that privileges language over rich audiovisual cues, striping away nuanced social cues; (3) evaluation practices reliant on static benchmarks, with socially grounded, human-centered assessments rare; and (4) Ethical discussions focused mainly on legal and rights-related risks (e.g., privacy), leaving societal risks (e.g., deception) overlooked—or at best acknowledged but left unaddressed. We outline a research agenda for evaluating socially competent, ethically informed, and interaction-aware multi-modal systems.
@@ 1. Introduction
This paper addresses the gap through a systematic literature review of LLM-powered multimodal systems that are designed to understand human behavior. Our analysis synthesizes research across technical venues in AI and applied communities such as HCI, education, and healthcare. Specifically, we examine the following four perspectives:
  • RQ1 (application): What is the application scope of LLM-powered multimodal systems for human behavior understanding in terms of target behaviors, interaction contexts, and operationalized social competencies?
  • RQ2 (technical): How are LLM-powered multimodal systems technically operationalized for human behavior understanding?
  • RQ3 (evaluative): What methodologies and metrics are employed to evaluate the performance of LLM-powered multimodal systems in understanding human behavior?
  • RQ4 (ethical): What are the primary ethical challenges and risks associated with developing and deploying LLM-powered multimodal systems for human behavior understanding?
@@ 3.3. Coding Procedure
4. RQ1: Scope and Nature of LLM-powered Multimodal Systems in Human Behavior Understanding
@@
4.1. Application Domains and Targeted Behavioral Cues
To map the landscape of LLM-powered multimodal systems in human behavior understanding, we analyzed both the application domains these systems target, and the types of human behaviors they are designed to recognize and interpret.
@@ 4.2. Behavior Characteristics
To characterize the nature of behaviors analyzed in LLM-powered multimodal systems, we examined three key dimensions: interaction structure (i.e., the number of interacting participants), temporal scale (i.e., the time span of behavior under analysis), and social units (i.e., the building blocks of social analysis). Together, these dimensions offer a structured view of the complexity of human behavior addressed by current systems (Figure. 4).
@@ 4.4. Operationalized Social Intelligence Competencies
Major Takeaways–Scope and Nature of Social Intelligence (RQ1) Dominance of Perception & Recognition: Nearly all of the 176 works focus on social perception (100%) and reasoning (95%), while social interaction (41.8%) and creativity (18%) are rarely implemented. Short-Term, Individual-Centric Focus: 94% of studies analyze immediate or short-term behavior and 63% focus on single-person analysis, with very limited modeling of long-term trajectories or multi-party dynamics. Uneven Cue Usage: Verbal & language cues dominate, while nuanced non-verbal signals like gaze, facial expressions, and prosody are underutilized despite their importance for social intelligence. Observer Role Prevails: Most systems act as post-hoc observers rather than active participants, highlighting a gap between analyzing and co-constructing social interaction.
5. RQ2: Technical Operationalization of LLM-powered Multimodal Systems in Human Behavior Understanding
@@ 5.2. Operationalization Technique
6. RQ3: Evaluation Methodologies and Metrics for LLM-powered Multimodal Systems in Human Behavior Understanding
@@ 6.2. System Evaluation Mechanism
7. RQ4: Ethical Challenges and Risks of LLM-powered Multimodal Systems in Human Behavior Understanding